Compound amount


2) (a) find the simple interest of 60,000 for one year at 10%

(b) what is the compound interest of the same investment is compounded semi annually ?

(c) how much is the compound interest greater than the simple interest ?


1
Expert's answer
2021-03-22T03:23:05-0400

(a) find the simple interest of 60,000 for one year at 10%


A = P ( 1+ r) T


A = 6000( 1 +0.10 ) 1

A = 6000(1.10)1

A = 66000

I = 66000- 60000= 6000


(b) what is the compound interest of the same investment is compounded semi annually ?

A = P ( 1+ 0.05)1/2

A = 60000( 1 +0.05 ) 1/2

A = 60000(1+0.05)1/2

A = 61482


Therefore:


A = P ( 1+ 0.05)1/2

A = 61482( 1 +0.05) 1/2

A = 61482(1.05)1/2

A = 63000

I = 63000-60000= 3000


(c) how much is the compound interest greater than the simple interest ?


63000-60000= 3000